What is the u substitution choice for "{\int \frac{\sin{\sqrt{x}}}{\sqrt{x}}dx}"


1) u=x

2) "u=\dfrac{1}{\sqrt{x}}"


3) u=x


4) None of the above



1
Expert's answer
2021-08-23T04:50:55-0400

u substitution should be

"u=\sqrt{x}"


to have "\frac{1}{2\sqrt{x}}dx=du"


Hence,

none of the above